an airplane files 500 miles with the wind for 2 hours and returns against the wind in 3 hours find the speed of the wind and the air speed of the plane
Answer by checkley77(12844)

DISTANCE=RATE*TIME
500/S=2
500=2S
S=500/2
S=250 MPH SPEED WITH THE WIND
500/S=3
3S=500
S=500/3
S=166.667 SPEED AGAINST THE WIND
(250-166.667)/2=416.667/2=208.333 IS THE SPEED OF THE PLANE.
THUS THE SPEED OF THE WIND IS:
250-208.333=41.667 MPH.
PROOF:
166.667+41.667=208.333 MPH.